A vehicle moving with a uniform acceleration of 2m/s has a velocity of 4m/s at a certain time. what will its velocity be?


1
Expert's answer
2019-10-22T10:39:27-0400

We can find the velocity of the vehicle from the kinematic equation:


"v = v_{0} + at,"

here, "v_0" is the initial velocity of the vehicle, "a" is the acceleration of the vehicle and "t" is the time.

Then, from this equation we can find the velocity of the vehicle 1 second later:


"v= 4 \dfrac{m}{s} + 2 \dfrac{m}{s^2} \cdot 1 s = 6 \dfrac{m}{s}."

Answer:

"v= 6 \dfrac{m}{s}."